OpenStudy (anonymous):

A professional baseball team has 9 picthers, 5 insiders, and 11 other players. if 3 players names are selected at ramdom, determine the prob that all 3 are infielders?

OpenStudy (anonymous):

You already did this problem...

OpenStudy (anonymous):

Just with different numbers

OpenStudy (anonymous):

how many total players are there?

OpenStudy (anonymous):

wouldnt it be 3/25?

OpenStudy (anonymous):

no, not so simple, because each time you select one, there is one less to select

OpenStudy (anonymous):

the first choice has a prob of 5/25 = 1/5

OpenStudy (anonymous):

the second has prob of 4/24 = 1/6

OpenStudy (anonymous):

RIGHT 5/25! not 3!

OpenStudy (anonymous):

and the third choice has a prob of 3/23 (doesn't reduce)

OpenStudy (anonymous):

so 1/5*1/6*3/23 = hmm... the 3 and 6 cancel, leaving a two in the denominator...

OpenStudy (anonymous):

1/(60*23) = 1/1380

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

1/230

OpenStudy (anonymous):

Wait im confused! o_o

OpenStudy (anonymous):

I meant 1/(5*2*23) instead

OpenStudy (anonymous):

Ohhhh ok

OpenStudy (anonymous):

ok I almost the same one now